(a) If the molar solubility of PbBr2 at 25 oC is 0.0118 mol/L, what is the Ksp at this temperature?

Ksp =

b) It is found that 6.25e-06 g of Cd3(PO4)2 dissolves per 100 mL of aqueous solution at 25 oC. Calculate the solubility-product constant for Cd3(PO4)2.

Ksp =

(c) The Ksp of Sc(OH)3 at 25 oC is 2.22e-31. What is the molar solubility of Sc(OH)3?

solubility = mol/L
Answer #1

a)

At equilibrium:

PbBr2 <----> Pb2+ + 2 Br-

   s 2s

Ksp = [Pb2+][Br-]^2

Ksp = (s)*(2s)^2

Ksp = 4(s)^3

Ksp = 4(1.18*10^-2)^3

Ksp = 6.572*10^-6

Answer: 6.57*10^-6

b)

S = 6.52*10^-6 g / 100 mL

S = 6.52*10^-6 g / 0.1 L

S = 6.52*10^-5 g /L

Molar mass of Cd3(PO4)2,

MM = 3*MM(Cd) + 2*MM(P) + 8*MM(O)

= 3*112.4 + 2*30.97 + 8*16.0

= 527.14 g/mol

Molar mass of Cd3(PO4)2= 527.14 g/mol

s = 6.52*10^-5 g/L

To covert it to mol/L, divide it by molar mass

s = 6.52*10^-5 g/L / 527.14 g/mol

s = 1.237*10^-7 mol/L

At equilibrium:

Cd3(PO4)2 <----> Cd2+ + 2 PO43-

   s 2s

Ksp = [Cd2+][PO43-]^2

Ksp = (s)*(2s)^2

Ksp = 4(s)^3

Ksp = 4(1.237*10^-7)^3

Ksp = 7.569*10^-21

Answer: 7.57*10^-21

C)

At equilibrium:

Sc(OH)3 <----> Sc3+ + 3 OH-

   s 3s

Ksp = [Sc3+][OH-]^3

2.22*10^-31=(s)*(3s)^3

2.22*10^-31= 27(s)^4

s = 9.522*10^-9 M

Answer: 9.52*10^-9 M